OpenStudy (anonymous):

find the following product. (y^2+7y-4) (y+6)

OpenStudy (anonymous):

FOIL it out.

OpenStudy (anonymous):

yes... carry out the distributive property twice: \(\large (y^2+7y-4) \color {red}{(y+6)} \) = \(\large y^2\cdot \color {red}{(y+6)}+7y\cdot \color {red}{(y+6)}-4\cdot \color {red}{(y+6)} \) from here carry out distributive property once more then combine like terms.....

OpenStudy (anonymous):

y^3 + 6y^2 + 7y^2 + 42y - 4y - 24, hope this helps alittle bit more. :]

OpenStudy (anonymous):

y^3+55y^5-28y

OpenStudy (anonymous):

You distributed and/or collected like terms incorrectly.

OpenStudy (anonymous):

y^3 + (6y^2 + 7y^2) + (42y - 4y) - 24 = y^3 + 13y^2 + 38y - 34 should have been your answer.

OpenStudy (anonymous):

what is the correct way to combine the like terms?

OpenStudy (anonymous):

the last term is -24.

OpenStudy (anonymous):

Like terms = Share a variable with the same exponents. For example y^2 + y^2 = 2y^2, or a^4b^2 + a^4b^2 = 2a^4b^2. And @qpHalcy0n yes. Sorry. Typo. :D

OpenStudy (anonymous):

As it turns out, 3 is very close to 2 :]

OpenStudy (anonymous):

Lol yes. Yes it is. D: